A particle moves along the x axis. Its position as a function of time is given by

x = 6 . 8 t + 8 . 5 t 2 , where t is in seconds and x is in meters. What is the acceleration as a

function of time?


1
Expert's answer
2021-04-05T11:09:42-0400

Let's first find the velocity of the particle:


"v=\dfrac{dx}{dt}=6.8+17t."

Finally, we can find the acceleration of the particle:


"a=\dfrac{dv}{dt}=17\ \dfrac{m}{s^2}."